OpenStudy (anonymous):

To make green paint, 3 parts yellow paint are mixed with 2 parts blue pain. Janis has 12L of blue paint. How much yellow paint does she need?

OpenStudy (anonymous):

Green paint= 3 parts yellow mixed with 2 parts blue ratio is 3:2 for yellow to blue x :12L solve for x \[\frac{x}{3}=\frac{12}{2}\]

OpenStudy (anonymous):

would x equal 6?

OpenStudy (anonymous):

you have to bring the 3 over.

OpenStudy (anonymous):

\[\frac{x}{3}=6\]